OpenStudy (dmezzullo):

To graph the equation y equals one half times x plus 3 using the slope and y-intercept, you would have to do which of the following? Graph the y-intercept, 3. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -3. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -3. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-i

OpenStudy (dmezzullo):

Graph the y-intercept, 3. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points.

OpenStudy (anonymous):

y=x/2 + 3 it is right?

OpenStudy (anonymous):

or y=(x+3)/2 ?

OpenStudy (dmezzullo):

idk thats y im askin u guys

OpenStudy (anonymous):

You need to be sure of the question first. "y equals one half times x plus 3" Write me the mathematical equation. Its part of the question.

OpenStudy (dmezzullo):

wut u mean?

OpenStudy (anonymous):

So its y=x/2 + 3. Hmm give me a min.

OpenStudy (dmezzullo):

okie dokie

OpenStudy (dmezzullo):

@nubeer

OpenStudy (anonymous):

I can tell you how to plot the line, but i dont think i know this method.

OpenStudy (stamp):

see attachment

OpenStudy (stamp):

whoops that x-int should be -6

OpenStudy (dmezzullo):

wait so which multiple choice answer should it be?

OpenStudy (dmezzullo):

@stamp

OpenStudy (dmezzullo):

??

OpenStudy (stamp):

I mean dude I am not going to tell you the answer, I graphed the equation for you. Based on the graph and work I have provided you, YOU should be able to answer YOUR question. I am here to help you, and doing everything for you does not equal helping you.

OpenStudy (dmezzullo):

Oh ok thanks

OpenStudy (dmezzullo):

(: